# Optical Injection and Detection of Long-Lived Interlayer Excitons in van der Waals Heterostructures

### v1

@@ Acknowledgements.
We thank Xiaobo Lu for fabricating Device 1. We thank Ivan Amelio, Haydn S. Adlong, Arthur Christianen and Igor Khanonkin for inspiring discussions. This work was supported by the Swiss National Science Foundation (SNSF) under Grant No. 200021-204076. A. M. S. acknowledges funding from an ETH Postdoctoral Fellowship. K.W. and T.T. acknowledge support from the JSPS KAKENHI (Grant Numbers 21H05233 and 23H02052), the CREST (JPMJCR24A5), JST and World Premier International Research Center Initiative (WPI), MEXT, Japan.References
@@
- Combescot et al. [2017] M. Combescot, R. Combescot, and F. Dubin, Reports on Progress in Physics 80, 066501 (2017).
@@
- Qi et al. [2023] R. Qi, A. Y. Joe, Z. Zhang, Y. Zeng, T. Zheng, Q. Feng, J. Xie, E. Regan, Z. Lu, T. Taniguchi, K. Watanabe, S. Tongay, M. F. Crommie, A. H. MacDonald, and F. Wang, Nat. Commun. 14, 8264 (2023).
@@
- Snoke [2011] D. Snoke, Advances in condensed matter physics 2011, 938609 (2011).
@@
- Kang et al. [2013] J. Kang, S. Tongay, J. Zhou, J. Li, and J. Wu, Appl. Phys. Lett. 102, 012111 (2013).
- Hattori et al. [2016] Y. Hattori, T. Taniguchi, K. Watanabe, and K. Nagashio, ACS Appl. Mater. Interfaces 8, 27877 (2016).
- Note [2] In contrast, MoSe2 displays a single AP peak under both electron and hole doping, making carrier-type identification more ambiguous. Additional confirmation comes from Device 3, a MoS2/WSe2 heterostructure, where the MoS2 layer clearly exhibits two AP features under electron doping, consistent with our interpretation of the spectra in Devices 1 and 2 (see Fig. S6).
@@
- Popert et al. [2022] A. Popert, Y. Shimazaki, M. Kroner, K. Watanabe, T. Taniguchi, A. Imamoğlu, and T. Smoleński, Nano Lett. 22, 7363 (2022).
- Kim et al. [2025] J. Kim, H. Dery, and D. Van Tuan, arXiv preprint arXiv:2505.06984 (2025).
- Smoleński et al. [2019] T. Smoleński, O. Cotlet, A. Popert, P. Back, Y. Shimazaki, P. Knüppel, N. Dietler, T. Taniguchi, K. Watanabe, M. Kroner, et al., Phys. Rev. Lett. 123, 097403 (2019).
- Note [3] The parabolic trend in the logarithmic plot in Fig. 4 is clearly visible; similar non-exponential time dependence has also been observed by Cutshall et al. [16] in a time-resolved PL measurement.
- Uto et al. [2024] T. Uto, B. Evrard, K. Watanabe, T. Taniguchi, M. Kroner, and A. İmamoğlu, Physical Review Letters 132, 056901 (2024).
- Khorana [1969] B. Khorana, Physical Review 185, 299 (1969).
- Pascucci et al. [2022] F. Pascucci, S. Conti, D. Neilson, J. Tempere, and A. Perali, Physical Review B 106, L220503 (2022).
- Biagioni et al. [2024] G. Biagioni, N. Antolini, B. Donelli, L. Pezzè, A. Smerzi, M. Fattori, A. Fioretti, C. Gabbanini, M. Inguscio, L. Tanzi, et al., Nature 629, 773 (2024).
- Rapaport et al. [2005] R. Rapaport, G. Chen, S. Simon, O. Mitrofanov, L. Pfeiffer, and P. Platzman, Physical Review B—Condensed Matter and Materials Physics 72, 075428 (2005).
- Gärtner et al. [2007] A. Gärtner, L. Prechtel, D. Schuh, A. Holleitner, and J. Kotthaus, Physical Review B—Condensed Matter and Materials Physics 76, 085304 (2007).

### v2

@@ Acknowledgements.
We thank Xiaobo Lu for fabricating Device 1. We thank Ivan Amelio, Haydn S. Adlong and Igor Khanonkin for inspiring discussions. This work was supported by the Swiss National Science Foundation (SNSF) under Grant No. 200021-204076. A.M.S. and A.C. acknowledge funding from an ETH Postdoctoral Fellowship. K.W. and T.T. acknowledge support from the JSPS KAKENHI (Grant Numbers 21H05233 and 23H02052), the CREST (JPMJCR24A5), JST and World Premier International Research Center Initiative (WPI), MEXT, Japan.References
@@
- Combescot et al. [2017] M. Combescot, R. Combescot, and F. Dubin, Rep. Prog. Phys. 80, 066501 (2017).
@@
- Qi et al. [2023] R. Qi, A. Y. Joe, Z. Zhang, Y. Zeng, T. Zheng, Q. Feng, J. Xie, E. Regan, Z. Lu, T. Taniguchi, K. Watanabe, S. Tongay, M. F. Crommie, A. H. MacDonald, and F. Wang, Nature Communications 14, 8264 (2023).
@@
- Snoke [2011] D. Snoke, Adv. Condens. Matter Phys. 2011, 938609 (2011).
@@
- Note [2] We emphasize that the two types of doping can be spectrally distinguished in WSe2 [35]: the presence of a single attractive polaron (AP) resonance in the WSe2 spectrum unequivocally indicates hole doping, since electron doping of WSe2 leads to two AP features associated with the singlet and triplet trions. In contrast, MoSe2 displays a single AP peak under both electron and hole doping. Additional confirmation comes from Device 3, a MoS2/WSe2 heterostructure, where the MoS2 layer clearly exhibits two AP features under electron doping, consistent with our interpretation of the spectra in Devices 1 and 2 (see Fig. S6).
@@
- Popert et al. [2022] A. Popert, Y. Shimazaki, M. Kroner, K. Watanabe, T. Taniguchi, A. Imamoğlu, and T. Smoleński, Nano Letters 22, 7363 (2022).
- Kim et al. [2025] J. Kim, H. Dery, and D. Van Tuan, Phys. Rev. B 112, L041301 (2025).
- [25] A. Christianen, A. M. Seiler, A. Tüğen, and A. İmamoğlu, in preparation .
- Kiper et al. [2025] N. Kiper, H. S. Adlong, A. Christianen, M. Kroner, K. Watanabe, T. Taniguchi, and A. İmamoğlu, Phys. Rev. X 15, 011049 (2025).
- Note [3] The parabolic trend in the logarithmic plot in Fig. 4 is clearly visible; similar nonexponential time dependence has also been observed by Cutshall et al. [16] in a time-resolved PL measurement.
- Uto et al. [2024] T. Uto, B. Evrard, K. Watanabe, T. Taniguchi, M. Kroner, and A. İmamoğlu, Phys. Rev. Lett. 132, 056901 (2024).
- Khorana [1969] B. Khorana, Phys. Rev. 185, 299 (1969).
- Lagoudakis et al. [2010] K. G. Lagoudakis, B. Pietka, M. Wouters, R. André, and B. Deveaud-Plédran, Phys. Rev. Lett. 105, 120403 (2010).
- Abbarchi et al. [2013] M. Abbarchi, A. Amo, V. Sala, D. Solnyshkov, H. Flayac, L. Ferrier, I. Sagnes, E. Galopin, A. Lemaître, G. Malpuech, et al., Nat. Phys. 9, 275 (2013).
- Rapaport et al. [2005] R. Rapaport, G. Chen, S. Simon, O. Mitrofanov, L. Pfeiffer, and P. Platzman, Phys. Rev. B 72, 075428 (2005).
- Gärtner et al. [2007] A. Gärtner, L. Prechtel, D. Schuh, A. Holleitner, and J. Kotthaus, Phys. Rev. B 76, 085304 (2007).